OpenStudy (anonymous):

Is the equation y = 5(0.5)x an example of growth or decay?

OpenStudy (anonymous):

This is growth I think..

OpenStudy (anonymous):

Yea i think so too

OpenStudy (anonymous):

Positive? can you explain to me?

OpenStudy (anonymous):

@zordoloom can you help?

OpenStudy (anonymous):

Well I'm actually not 100% sure, we can represent 5 as a and (0.5)x and b

OpenStudy (anonymous):

This would be decay.

OpenStudy (anonymous):

basically in these equations b is between 0 and 1 then it is a decay, and if it is more then 1 it is a growth

OpenStudy (anonymous):

can you explain zordoloom?

OpenStudy (anonymous):

Just correct your equation. y = 5(0.5)x is not correct. It should be y = 5(0.5)^x

OpenStudy (anonymous):

Is this the correct equation? y = 5(0.5)^x?

OpenStudy (anonymous):

IS the x exponential?

OpenStudy (anonymous):

Yes